OpenStudy (anonymous):

Solve tan x < 2 sin x for [0,2pi)

OpenStudy (anonymous):

tan x < 2 sin x (sin x / cos x) < 2 sin x sin x < 2 sin x cos x sin x < sin 2x What next?

OpenStudy (anonymous):

i dont think you can multiply by cosx as it can be negative

OpenStudy (anonymous):

cosine is not positive on \((0,2\pi)\) so you cannot multiply both sides by cosine

OpenStudy (mertsj):

\[\frac{\sin x}{\cos x}-2\sin x <0\] \[\sin x(\frac{1}{\cos x}-2)<0\]

OpenStudy (anonymous):

\[tanx<2sinx\] \[tanx - 2sinx < 0\] \[\frac{(sinx - 2sinxcosx)}{cosx} < 0\] now consider critical values where denominator, numerator = 0 we have to consider sinx(1- 2cosx) = 0 cosx = 0 so get these values of x, then find the sign of \[\frac{(sinx - 2sinxcosx)}{cosx}\] between them. this will solve your inequality.

OpenStudy (anonymous):

have to make it positive or negative. what mertsj said

OpenStudy (anonymous):

alternatively just solve tanx = 2sinx then find the sign of tanx - 2sinx

OpenStudy (anonymous):